An inclined plane is 3.5 m long and rises o.95 meters. The weight of the object on the inclined plane is 9.6 N. Find a) angle of the inclined plane ? , b) What is the normal force ? , and c) The force to move the object.



1
Expert's answer
2021-04-11T20:06:21-0400

a) The angle of the inclined plane is


"\theta=\arcsin\frac{0.95}{3.5}=15.7º."


b) The normal force is


"N=W\cos\theta=9.6\cos15.7º=9.24\text{ N}."


c) The force to move the object

To move the object up the incline at constant velocity, the following force is required:


"F=W\sin\theta=9.6\sin15.7º=2.60\text{ N}."